how do you solve the equation

b+two thirds=2

Answer by nica_021794(8)

b+two thirds=2
b+2/3=2
Multiply the whole equation by the LCD:
3b+2=6
Transpose 2 to the right side of the equation:
3b=6-2
3b=4
Divide the whole equation by 3
(3b=4)/3
b=4/3